OpenStudy (anonymous):

Solve. x^2 – 10x = 5

OpenStudy (australopithecus):

x^(2) - 10x - 5 = 0 Now use the quadratic formula https://www.youtube.com/watch?v=EeVqtpuMFOU

OpenStudy (cwrw238):

or:- (x - 5)^2 - 25 = 5 x - 5 = +/- sqrt 30 thats the completing the square method